Showing posts with label Binomial. Show all posts
Showing posts with label Binomial. Show all posts

Understanding Binomial Distribution with Excel 2007

In this post we will try to understand how the Excel built-in Binomial distribution function can be used for calculating Binomial probabilities. Traditionally a binomial random variable is described by the following formula: 
$$b(x,n,p)=\binom{n}{x}p^x q^{n-x}$$

Where $p$ is the probability of success and $q$ is the probability of failure and can be related by the relation $p+q=1$, $n$ is the number of trails and $x$ is the value at which we want to evaluate the binomial probability .

The Excel 2007 has a built-in function of BINOMDIST() that takes arguments of $x,n,p$ and logical arguments for cumulative or point density estimate and return the value of Binomial Variable.

The Syntex of the function is :BINOMDIST(number_s,trials,probability_s,cum) 

Now we will see how the calculation that are done manually can be skipped if you use BINOMDIST() function, lets consider this example:

Find the probability of obtaining exactly three 2's if an ordinary dice is tossed 5 times.

Solution: The probability of success is this case is $p=\frac{1}{6},n=5,x=3$ gives:
$$b(3,5,\frac{1}{6})=\binom{5}{3}(\frac{1}{6})^{3}(\frac{5}{6})^{2}=\frac{5!}{3!2!}\cdot\frac{5!}{6!} = 0.032$$
Using Excel: The same can be calculated by  the formula: 

=Binomdist(3,5,0.1667,False)=0.032

Example # 02: The probabilities that a patient recovers from a rare blood disease is 0.4. if 15 people are known to have contracted this disease, what is the chance that (a) at lease 10 will survive (b) from 3 to 8 will survive and (c) exactly 5 will survive?

Solution: 

(a) For Probability $that at lease 10 will survive$:

$$P(x\geq 10) = 1-P(x\leq 10) = 1 - \sum_{x=0}^{9}b(x,15,0.4) = 1-0.9662 = 0.0338$$
(b) For Probability $that from 3 to 8 will survive$
$$P(3\leq x\leq8) = \sum_{x=3}^{8}b(x,15,0.4) = \sum_{x=0}^{8}b(x,15,0.4)-\sum_{x=2}^{8}b(x,15,0.4)$$
$$=0.9050-0.0271 = 0.8779$$

(c) For Probability $that exactly 5 will survive$
$$P(x = 5) = b(x,15,0.4) = \sum_{x=0}^{8}b(x,15,0.4)-\sum_{x=0}^{4}b(x,15,0.4)$$
$$=0.4032-0.02173 = 0.1859$$

Using Excel:

(a) =1-Binomdist(9,15,0.4,TRUE) = 0.0338
(b) =Binomdist(8,15,0.4,TRUE)-Binomdist(2,15,0.4,TRUE) = 0.8779
(c) =Binomdist(5,15,0.4,TRUE)-Binomdist(4,15,0.4,TRUE) = 0.1859

hence we can see how easily we can estimate the binomial probabilities using Excel 2007, I hope that you will like this post. Please give your feedback to improve the blog.

The Google Sheet Assignment - Splitting and Transposing Cell based on Delimiter - Part 1

I recently had opportunity to workout a formula with google sheets, where i realized how powerful google sheet formulas are. The problem ...